# Reconciling aging and slip state evolutions from laboratory-derived canons of rate-and-state frictionF

### v2

@@ 3.2 Transient evolution in ideal VS tests
In ideal VS tests, the interface slip rate $V$ is abruptly changed from one value ($V_{\rm before}$) to another ($V_{\rm after}$), following steady-state sliding at $V=V_{\rm before}$, which sets the strength to $\Phi(V_{\rm before})$ as the initial condition for the post-step evolution. In laboratory experiments, such a perfect step change of $V$ is approximately realized by abruptly changing the load-point velocity $V_{\rm m}$ from $V_{\rm before}$ to $V_{\rm after}$ on a sufficiently stiff loading apparatus. Following the velocity step, the strength $\Phi$ evolves from the old steady-state value $\Phi_{\rm SS}(V_{\rm before})$ to the new value $\Phi_{\rm SS}(V_{\rm after})$, exponentially with the slip displacement $\delta$ since the moment of the step. A critical, well-established experimental observation is that the characteristic slip distance $D_{\rm c}$ in this exponential evolution does not depend on either $V_{\rm before}$ or $V_{\rm after}$ (Ruina, 1980; Bhattacharya et al., 2015):
@@
The VS canon, idealized by using the infinite machine stiffness, clearly indicates the time-evolution law that governs the slip weakening and strengthening. We now derive that time-evolution law in a time-derivative form, which applies to whatever $V(t)$, including the real-word VS under finite stiffness, where non-Heaviside $V(t)$ arises from the elastic interaction and the RSF constitutive law (eq. 10), reproducing the non-exponential behaviors of the frictional stress under actual VS tests.
@@ 3.3 SHS tests under the negligible slip condition
If the interface slip rate $V$ during hold is strictly zero as intended, the observed log-t healing is entirely ascribed to the time-dependent change in strength (e.g., Rabinowicz, 1958; Dieterich, 1972, 1978). Within the TDF framework (eq. 22), a positive $F_{t}$ term expresses such purely time-dependent evolution as in the aging law (e.g., Dieterich, 1979; Ruina, 1983). However, in the majority of SHS tests, slow slip continues during the hold phase (see §3.3.1 for detailed explanation) to an extent sufficient for the slip law, which lacks the purely time-dependent evolution (i.e., $F_{t}=0$), to reproduce the observed log-t healing as a consequence of slip-dependent evolution expressed by the $F_{\delta}V$ term (Ruina, 1983; Beeler et al., 1994). Thus, the log-t healing observed in such SHS tests can be explained by either the slip law (as ‘apparent’ healing) or the aging law (as ‘purely time-dependent’ healing). Although the behaviors predicted by the two laws are not the same even for such SHS tests (e.g., Beeler et al., 1994; Bhattacharya et al., 2017), the difference is not significant enough to confidently reject either laws (e.g., Bhattacharya et al., 2017), given the limited experimental accuracy and coverage of conditions. In contrast, as detailed in §3.3.1, some SHS tests were done at virtually zero slip rates during hold (Nakatani & Mochizuki, 1996) or at slip rates way too low for the slip law to reproduce the log-t healing. Nevertheless, significant log-t healing was observed in these tests. For example, in the experiments of Nakatani & Mochizuki (1996), shear stress during hold was kept as low as approximately $0.02\sigma$, reduced by approximately $0.85\sigma$ from the preceding steady-state sliding. The interface slip rate $V$ during the hold expected from the RSF constitutive law (eq. 8) is so low that slip during hold does not even reach the atomic spacing in rock-forming minerals within the lifetime of human experimenters. An absurdly huge $F_{\delta}$ would be required to ascribe the observed healing to slip-driven evolution under such low $V$. Hence, their observation of significant log-t healing implies that purely time-dependent healing does operate, i.e., $F_{t}>0$, at least in the holds at virtually zero slip rates.
@@ 3.3.1 Technical details of SHS tests
Meanwhile, Nagata et al. (2008) have demonstrated that acoustic transmissivity across the frictional interface is linearly related to $\Phi$ (observable as per eq. 10). With this method, $\Phi_{\rm end}$ or even the continuous record of $\Delta\Phi(t)$ throughout the hold period can be monitored in-situ, whereby the reloading operation and the above-mentioned correction become unnecessary. Although the linear relationship between acoustic transmissivity and $\Phi$ breaks during reloading to $\tau_{\rm peak}$ (e.g., Nagata et al. 2008), which is the initial part of slip-weakening of $\Phi$, the present paper refers to the acoustic-transmissivity- $\Phi$ linearity only for SS and hold phase of SHS tests. Furthermore, for transparent materials, optical transmissivity can play the same role (Dieterich & Kilgore, 1994; Nagata et al., 2014). It also supports the ATF (Bowden & Tabor, 1964) interpretation of $\Phi$ because optical transmissivity is proportional to the real contact area. These acoustic/optical results, though indirect measurements of the strength $\Phi$, may also be usable as a basis to formulate the canonical log-t healing behavior.
@@ 3.3.2 Canonical behavior and evolution-law constraint
For evolution laws having $f(V,\theta)$ form, including the aging law and the slip law (eqs. 4 and 5), it may not be appropriate to regard eq. (49) as a requirement. As long as we presume that $\dot{\theta}$ only depends on the instantaneous $V$ and $\theta$, namely $\dot{\theta}=f(V,\theta)$, eq. (49) follows three of the other experimental requirements, eqs. (26), (48), and (67). See Appendix D for proof. In that sense, eq. (49) is a theoretical consequence of other requirements rather than an additional requirement. Nonetheless, to keep the generality of our derivation, we position eq. (49) as a requirement from SHS tests, which would not harm the present development.
@@
Equation (62) was used in the above derivation. We note that eq. (66) can also be obtained from eq. (60) via the same variable conversion (eq. 15). The $c_{*}$ in eq. (60) disappears after this variable conversion, indicating that the $c_{*}$-indefiniteness in the $\theta$’s evolution law (eq. 60) is merely apparent.
@@
Nakatani & Mochizuki (1996) did constant-$\tau$ SHS tests at various levels of $\tau_{\rm hold}$, ranging from $0.02\sigma$ to $0.084\sigma$, approximately corresponding to $\Delta\tau$ of $0.83\sigma$ to $0.01\sigma$. With $A=0.01\sigma$, as typical for nominally bare granite interfaces they used, the corresponding range of $\Delta\tau/A$ is 1–83. Their healing data (Figure 4a of Nakatani & Mochizuki, 1996) indicate that $t_{\rm c}$ was about 3–30 ms, with no evident systematic dependence on $\tau_{\rm hold}$. For this range of $\Delta\tau/A$, the slip law predicts that $t_{\rm c}$ increases with $\Delta\tau$ by a factor of about $10^{34}$. Even if we exclude tests with $\tau_{\rm hold}<0.3\sigma$ (i.e., $\Delta\tau>0.55\sigma$), where their $B_{\rm heal}$ showed a systematic dependence on $\tau_{\rm hold}$, the slip law still predicts a variation in $t_{\rm c}$ by a factor of about $10^{22}$. Thus, the slip law’s prediction on $t_{\rm c}$ strongly contradicts the laboratory observation of $t_{\rm c}$ little affected by $\tau_{\rm hold}$ over its entire range from zero to $\tau_{\rm SS}(V_{\rm prior})$.
We can also argue against the slip law based on the absolute value of $t_{\rm c}$ predicted by the law. The slip law’s prediction of $t_{\rm c}$ (eq. A.26) can be rewritten into: $t_{\rm c|slip}=(D_{\rm c}/V_{\rm prior})\times(A/\Delta\tau)\times(A/B)\times\exp(-\Delta\tau/A)$. [Note that $D_{\rm c}/V_{\rm prior}$, instead of $L/V_{\rm prior}$, appears as the characteristic time of the apparent log-t healing predicted by the slip law, or more generally, by any apparent healing deriving from the same form of ODE ($\dot{\theta}=-(L/D_{\rm c})\Omega\ln\Omega$). This is because the governing ODE can be written as $\dot{\theta}^{\prime}=-(\theta^{\prime}V/D_{\rm c})\ln(\theta^{\prime}V/D_{\rm c})$ using a state variable $\theta^{\prime}$ whose steady-state value is $D_{\rm c}/V$.] According to the slip law, $D_{\rm c}$ can be directly observed in VS tests (Fig. 2a). Assuming the typical value of $D_{\rm c}=1\mu$m, and $V_{\rm prior}$ of about $20\mu$m/s in their tests, $D_{\rm c}/V_{\rm prior}=0.1$s. For typical interfaces with $A\simeq B$, $t_{\rm c|slip}$ is about a century for tests with $\Delta\tau=27A$. For their lowest $\tau_{\rm hold}$ of $0.02\sigma$, (i.e., $\Delta\tau=83A$), $t_{\rm c|slip}$ predicted by eq. (A.26) is about $10^{25}$ years, $10^{15}$ times longer than the age of our observable universe, whereas those experimenters found steadfast log-t healing (about $0.01\sigma$ strengthening per 10-fold increase in $t_{\rm h}$) for the much shorter range of $t_{\rm h}$ (30–3000s) they tested.
@@
Note, however, the above argument of Nakatani & Scholz (2006) presumed the NSC, which may not have been the case in the stationary-loadpoint hold phase of SHS tests by Marone (1998a) they analyze. Although we think that $L\sim D_{\rm c}$ is likely, it is not as well-founded as the three canons are. We first assume $L\sim D_{\rm c}$ (eq. 70) and utilize it as an auxiliary constraint in developing evolution laws in §4.3 and then consider the effects of $L/D_{\rm c}$ significantly differing from unity in §4.4.
@@ 4 Development of new evolution law
Before proceeding to §4.1, we here point out that eq. (73) is, in fact, the aging law under the NSC. This fact can be proven by showing that, for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, the NSC is possible only in a narrow range of $\Omega$. Given eqs. (72) and (73), $|f_{\delta}V|\ll|f_{t}|$ if and only if $\Omega\sim 1\cup\Omega\ll D_{\rm c}/L$. [Strictly speaking, for $|\ln(D_{\rm c}/L)|\gg 1$, the necessary and sufficient condition for $|f_{\delta}V|\ll f_{t}$ is $\Omega\sim 1\cup\Omega\ll(D_{\rm c}/L)/|\ln(D_{\rm c}/L)|$. However, unless $D_{\rm c}$ and $L$ differ by very many orders, this difference is negligible, so that basically the former conditions suitable when $|\ln(D_{\rm c}/L)|\lesssim 1$ shall be considered hereafter.] Moreover, as VS tests involve situations where $\Omega\sim\Omega_{\rm SS}=1$ (eq. 71), the VS constraint eq. (44) precludes the possibility that $\Omega\sim\Omega_{\rm SS}$ satisfies the NSC. Therefore, these conclude that, under eqs. (71)–(73) and (44), the NSC can exist only for $\Omega\ll D_{\rm c}/L$. Especially, $\Omega\to 0$ corresponds to the ideal SHS||||NSC tests, where $\tau_{\rm hold}$ is kept much lower than $\tau_{\rm SS}(V_{\rm prior})$ so that ideal stationary contact is almost achieved. As we have adopted the SHS canon (eq. 48) from such tests in §3.3, we regard $\Omega\to 0$ as the ideal NSC limit. For the aging law (eq. 4), where $D=D_{\rm c}=L$ is assumed, the above proposition (NSC $\subset\Omega\ll D_{\rm c}/L$) reduces to NSC $\subset\Omega\ll 1$, resulting in that eq. (73) is the aging law under the NSC. Thus, the reconciliation of the aging law and the slip law is equivalent to the reconciliation of the time-driven healing of (static) friction in SHS||||NSC tests and the slip-driven evolution of (dynamic) friction in VS tests.
@@
To reconcile the canons, we treat $B$, $L$, and $D_{\rm c}$ as constants. According to Nakatani & Mochizuki (1996) and Ryan et al. (2018), $B$ and $L$ do not change their orders for arbitrary $\tau=\tau_{\rm hold}>0$, and the $L\sim D_{\rm c}$ case is our main concern. Thus, we can neglect the variations in $B$, $L$, and $D_{\rm c}$ as second-order effects irrelevant to reconciling the two laws, although in applying the developed law, we will need to account for the $\tau$-dependence of $B$, which arises for $\tau_{\rm hold}<0.4\sigma$, where $B_{\rm heal}$ varies by factors; that dependence is a part of thermomechanical properties of the RSF parameters, typified by the temperature dependence of $A$ and $B$.

### v3

@@ 3.2 Transient evolution in ideal VS tests
In ideal VS tests, the interface slip rate $V$ is abruptly changed from one value ($V_{\rm before}$) to another ($V_{\rm after}$), following steady-state sliding at $V=V_{\rm before}$, which sets the strength to $\Phi_{\rm SS}(V_{\rm before})$ as the initial condition for the post-step evolution. In laboratory experiments, such a perfect step change of $V$ is approximately realized by abruptly changing the load-point velocity $V_{\rm m}$ from $V_{\rm before}$ to $V_{\rm after}$ on a sufficiently stiff loading apparatus. Following the velocity step, the strength $\Phi$ evolves from the old steady-state value $\Phi_{\rm SS}(V_{\rm before})$ to the new value $\Phi_{\rm SS}(V_{\rm after})$, exponentially with the slip displacement $\delta$ since the moment of the step. A critical, well-established experimental observation is that the characteristic slip distance $D_{\rm c}$ in this exponential evolution does not depend on either $V_{\rm before}$ or $V_{\rm after}$ (Ruina, 1980; Bhattacharya et al., 2015):
@@
The VS canon, idealized by using the infinite machine stiffness, clearly indicates the time-evolution law that governs the slip weakening and strengthening. We now derive that time-evolution law in a time-derivative form, which applies to whatever $V(t)$, including the real-world VS under finite stiffness, where non-Heaviside $V(t)$ arises from the elastic interaction and the RSF constitutive law (eq. 10), reproducing the non-exponential behaviors of the frictional stress under actual VS tests.
@@ 3.3 SHS tests under the negligible slip condition
If the interface slip rate $V$ during hold is strictly zero as intended, the observed log-t healing is entirely ascribed to the time-dependent change in strength (e.g., Rabinowicz, 1958; Dieterich, 1972, 1978). Within the TDF framework (eq. 22), a positive $F_{t}$ term expresses such purely time-dependent evolution as in the aging law (e.g., Dieterich, 1979; Ruina, 1983). However, in the majority of SHS tests, slow slip continues during the hold phase (see §3.3.1 for detailed explanation) to an extent sufficient for the slip law, which lacks the purely time-dependent evolution (i.e., $F_{t}=0$), to reproduce the observed log-t healing as a consequence of slip-dependent evolution expressed by the $F_{\delta}V$ term (Ruina, 1983; Beeler et al., 1994). Thus, the log-t healing observed in such SHS tests can be explained by either the slip law (as ‘apparent’ healing) or the aging law (as ‘purely time-dependent’ healing). Although the behaviors predicted by the two laws are not the same even for such SHS tests (e.g., Beeler et al., 1994; Bhattacharya et al., 2017), the difference is not significant enough to confidently decide on either laws (e.g., Bhattacharya et al., 2017), given the limited experimental accuracy and coverage of conditions. In contrast, as detailed in §3.3.1, some SHS tests were done at virtually zero slip rates during hold (Nakatani & Mochizuki, 1996) or at slip rates way too low for the slip law to reproduce the log-t healing. Nevertheless, significant log-t healing was observed in these tests. For example, in the experiments of Nakatani & Mochizuki (1996), shear stress during hold was kept as low as approximately $0.02\sigma$, reduced by approximately $0.85\sigma$ from the preceding steady-state sliding. The interface slip rate $V$ during the hold expected from the RSF constitutive law (eq. 8) is so low that slip during hold does not even reach the atomic spacing in rock-forming minerals within the lifetime of human experimenters. An absurdly huge $F_{\delta}$ would be required to ascribe the observed healing to slip-driven evolution under such low $V$. Hence, their observation of significant log-t healing implies that purely time-dependent healing does operate, i.e., $F_{t}>0$, at least in the holds at virtually zero slip rates.
@@ 3.3.1 Technical details of SHS tests
Meanwhile, Nagata et al. (2008) have demonstrated that acoustic transmissivity across the frictional interface is linearly related to $\Phi$ (observable as per eq. 10). With this method, $\Phi_{\rm end}$ or even the continuous record of $\Delta\Phi(t)$ throughout the hold period can be monitored in-situ, whereby the reloading operation and the above-mentioned correction become unnecessary. Although the linear relationship between acoustic transmissivity and $\Phi$ breaks during reloading to $\tau_{\rm peak}$ (e.g., Nagata et al., 2008), which is the initial part of slip-weakening of $\Phi$, the present paper refers to the acoustic-transmissivity- $\Phi$ linearity only for SS and hold phase of SHS tests. Furthermore, for transparent materials, optical transmissivity can play the same role (Dieterich & Kilgore, 1994; Nagata et al., 2014). It also supports the ATF (Bowden & Tabor, 1964) interpretation of $\Phi$ because optical transmissivity is proportional to the real contact area. These acoustic/optical results, though indirect measurements of the strength $\Phi$, may also be usable as a basis to formulate the canonical log-t healing behavior.
@@ 3.3.2 Canonical behavior and evolution-law constraint
However, for evolution laws having $f(V,\theta)$ form, including the aging law and the slip law (eqs. 4 and 5), it may not be appropriate to regard eq. (49) as a part of the canons. As long as we presume that $\dot{\theta}$ only depends on the instantaneous $V$ and $\theta$, namely $\dot{\theta}=f(V,\theta)$, eq. (49) follows three of the other empirical formulae, eqs. (26), (48), and (67), as shown in Appendix D. In this sense, eq. (49) is a theoretical consequence of other empirical formulae rather than a requirement. Nonetheless, for the sake of prudence, we position eq. (49) as a requirement from SHS tests, which would not harm the present development.
@@
Equation (62) was used in the above derivation. We note that eq. (66) can also be obtained from eq. (60) via the same variable conversion (eq. 15). The $c_{*}$ in eq. (60) disappears after this variable conversion without choosing specific $c_{*}$ values, indicating that the $c_{*}$-indefiniteness in the $\theta$’s evolution law (eq. 60) is merely apparent, not existent in $\Phi$, the observable.
@@
Nakatani & Mochizuki (1996) did constant-$\tau$ SHS tests at various levels of $\tau_{\rm hold}$, ranging from $0.02\sigma$ to $0.84\sigma$, approximately corresponding to $\Delta\tau$ of $0.83\sigma$ to $0.01\sigma$. With $A=0.01\sigma$, as typical for nominally bare granite interfaces they used, the corresponding range of $\Delta\tau/A$ is 1–83. Their healing data (Figure 4a of Nakatani & Mochizuki, 1996) indicate that $t_{\rm c}$ was about 3–30 ms, with no evident systematic dependence on $\tau_{\rm hold}$. For this range of $\Delta\tau/A$, the slip law predicts that $t_{\rm c}$ increases with $\Delta\tau$ by a factor of about $10^{34}$. Even if we exclude tests with $\tau_{\rm hold}<0.3\sigma$ (i.e., $\Delta\tau>0.55\sigma$), where their $B_{\rm heal}$ showed a systematic dependence on $\tau_{\rm hold}$, the slip law still predicts a variation in $t_{\rm c}$ by a factor of about $10^{22}$. Thus, the slip law’s prediction on $t_{\rm c}$ strongly contradicts the laboratory observation of $t_{\rm c}$ little affected by $\tau_{\rm hold}$ over its entire range from zero to $\tau_{\rm SS}(V_{\rm prior})$.
We can also argue against the slip law based on the absolute value of $t_{\rm c}$ predicted by the law. The slip law’s prediction of $t_{\rm c}$ (eq. A.26) can be rewritten into: $t_{\rm c|slip}=(D_{\rm c}/V_{\rm prior})\times(A/\Delta\tau)\times(A/B)\times\exp(-\Delta\tau/A)$. [Note that $D_{\rm c}/V_{\rm prior}$, instead of $L/V_{\rm prior}$, appears as the characteristic time of the apparent log-t healing predicted by the slip law, or more generally, by any apparent healing deriving from the same form of ODE ($\dot{\theta}=-(L/D_{\rm c})\Omega\ln\Omega$). This is because the governing ODE can be written as $\dot{\theta}^{\prime}=-(\theta^{\prime}V/D_{\rm c})\ln(\theta^{\prime}V/D_{\rm c})$ using a state variable $\theta^{\prime}$ whose steady-state value is $D_{\rm c}/V$.] According to the slip law, $D_{\rm c}$ can be directly observed in VS tests (Fig. 2a). Assuming the typical value of $D_{\rm c}=1$ $\mu$m, and $V_{\rm prior}$ of about $20$ $\mu$m/s in their tests, $D_{\rm c}/V_{\rm prior}=50$ ms. For typical interfaces with $A\simeq B$, $t_{\rm c|slip}$ is about a century for tests with $\Delta\tau=27A$. For the lowest $\tau_{\rm hold}$ of $0.02\sigma$ in Nakatani & Mochizuki (1996), $\Delta\tau=83A$, and $t_{\rm c|slip}$ predicted by eq. (A.26) is about $10^{25}$ years, $10^{15}$ times longer than the age of our observable universe, whereas those experimenters found steadfast log-t healing (about $0.01\sigma$ strengthening per 10-fold increase in $t_{\rm h}$) for the much shorter range of $t_{\rm h}$ (30–3000 s) they tested.
@@
Note, however, the above argument of Nakatani & Scholz (2006) presumed the NSC, which may not have been the case in the stationary-loadpoint hold phase of SHS tests by Marone (1998a) they analyze. As shown in Appendix B, stationary-loadpoint SHS has difficulty in probing the NSC regime. Rather, support for $t_{\rm c}=L_{\rm heal}/V_{\rm prior}$ (eq. 67) and thus $L\sim D_{\rm c}$ (eq. 70) comes from the constant-$\tau_{\rm hold}$ SHS tests of Nakatani & Mochizuki (1996), where the same order of $t_{\rm c}$ values were found in tests with low $\tau_{\rm hold}$ satisfying the NSC and tests with so high $\tau_{\rm hold}$ that cannot satisfy the NSC; the SHS||||NSC and SS canons (i.e., the aging law under the NSC) predict $t_{\rm c}=L_{\rm heal}/V_{\rm prior}$ under the NSC ($|f_{t}|\gg|f_{\delta}V|$), whereas the VS and SS canons (i.e., the slip law) assert $t_{\rm c}\sim D_{\rm c}/V_{\rm prior}$ when $|f_{t}|\ll|f_{\delta}V|$ (Fig. 5). Although we think that eqs. (67) and (70) are likely, they are not as well-founded as thought to be. Nonetheless, we proceed by assuming eq. (67) as it is tied to the presumption of $B=B_{\rm heal}$ (Appendix B). Meanwhile, regarding eq. (70), although we first assume $L\sim D_{\rm c}$ (eq. 70) and utilize it as an auxiliary constraint in developing evolution laws in §4.3, we consider the effects of $L/D_{\rm c}$ significantly differing from unity in §4.4.
@@ 4 Development of new evolution law
Before proceeding to §4.1, we here point out that eq. (73) is, in fact, the aging law under the NSC. This fact can be proven by showing that, for $\theta,\theta_{\rm SS}\gg\theta_{\rm X}$, the NSC is possible only in a narrow range of $\Omega$. Given eqs. (72) and (73), $|f_{\delta}V|\ll|f_{t}|$ if and only if $\Omega\sim 1\cup\Omega\ll D_{\rm c}/L$. [Strictly speaking, for $|\ln(D_{\rm c}/L)|\gg 1$, the necessary and sufficient condition for $|f_{\delta}V|\ll f_{t}$ is $\Omega\sim 1\cup\Omega\ll(D_{\rm c}/L)/|\ln(D_{\rm c}/L)|$. However, unless $D_{\rm c}$ and $L$ differ by very many orders, this difference is negligible, so that basically the former conditions suitable when $|\ln(D_{\rm c}/L)|\lesssim 1$ shall be considered hereafter.] Moreover, as VS tests involve situations where $\Omega\sim\Omega_{\rm SS}=1$ (eq. 71), the VS constraint eq. (44) precludes the possibility that $\Omega\sim\Omega_{\rm SS}$ satisfies the NSC. Therefore, these conclude that, under eqs. (71)–(73) and (44), the NSC can exist only for $\Omega\ll D_{\rm c}/L$. Especially, $\Omega\to 0$ corresponds to the ideal SHS||||NSC tests, where $\tau_{\rm hold}$ is kept much lower than $\tau_{\rm SS}(V_{\rm prior})$ so that ideal stationary contact is almost achieved. As we have adopted the SHS canon (eqs. 48 and 49) from such tests in §3.3, we regard $\Omega\to 0$ as the ideal NSC limit. For the aging law (eq. 4), where $D=D_{\rm c}=L$ is assumed, the above proposition (NSC $\subset\Omega\ll D_{\rm c}/L$) reduces to NSC $\subset\Omega\ll 1$, resulting in that eq. (73) is the aging law under the NSC. Thus, the reconciliation of the aging law and the slip law is equivalent to the reconciliation of the time-driven healing of (static) friction in SHS||||NSC tests and the slip-driven evolution of (dynamic) friction in VS tests.
@@
To reconcile the canons, we treat $B$, $L$, and $D_{\rm c}$ as constants. According to Nakatani & Mochizuki (1996) and Ryan et al. (2018), $B$ and $L$ do not change their orders for arbitrary $\tau=\tau_{\rm hold}>0$, and the $L\sim D_{\rm c}$ case is our main concern. Thus, we can neglect the variations in $B$, $L$, and $D_{\rm c}$ as secondary effects irrelevant to reconciling the two laws, although in applying the developed law, we will need to account for the $\tau$-dependence of $B$, which arises for $\tau_{\rm hold}<0.4\sigma$, where $B_{\rm heal}$ varies by factors; that dependence is a part of thermomechanical properties of the RSF parameters, typified by the temperature dependence of $A$ and $B$.